The length of a rectangle is 5m longer than its width.
If the perimeter of the rectangle is 38m , find its area.









Answer by algebrahouse.com(1659)

"The length of a rectangle is 5m longer than its width.
If the perimeter of the rectangle is 38m , find its area."

x = width
x + 5 = length

2x + 2(x + 5) = 38 {perimeter equals two times length plus two times width}
2x + 2x + 10 = 38 {used distributive property}
4x + 10 = 38 {combined like terms}
4x = 28 {subtracted 10 from both sides}
x = 7 {divided both sides by 4}
x + 5 = 12 {substituted 7, in for x, into x + 5}

width = 7m
length = 12m
area = 84 m^2 {area = length x width}
www.algebrahouse.com